OpenStudy (anonymous):

let f(x)=x-2 and g(x)=x^2-7x-9. find f(g(-1))

OpenStudy (anonymous):

are you still looking for the inverse?

OpenStudy (anonymous):

wait i put it wrong let me try again

OpenStudy (anonymous):

let f(x)=x^2-7x-9. find f(g(-1))

OpenStudy (anonymous):

and yea im looking for the same thing i was last time

OpenStudy (anonymous):

i don't think so this time i think this is \[f(g(-1))\]

OpenStudy (anonymous):

but i could be wrong could you take a screen shot of the problem?

OpenStudy (anonymous):

your right thats wat it is and my computer dosnt have screen shot im sorry

OpenStudy (anonymous):

if it is \(f(g(-1))\) your first step is to compute \(g(-1)\) do you know how to do that?

OpenStudy (anonymous):

make the -1 dissapear

OpenStudy (anonymous):

substitute -1 in g(x) after you find value of g(x), substitute that value to f(x)

OpenStudy (anonymous):

lol yeah, make it go away

OpenStudy (anonymous):

i got 21 for my answer

OpenStudy (anonymous):

\[ g(x)=x^2-7x-9\]\[ g(\heartsuit)=\heartsuit^2-7\heartsuit-9\]\[ g(-1)=(-1)^2-7(-1)-9\]

OpenStudy (anonymous):

so the question is, what is \[(-1)^2-7\times (-1)-9\]

OpenStudy (anonymous):

I want to confirm something. Is g(x)=x^2-7x-9 and f(x)=x^2-7x-9?

OpenStudy (anonymous):

@eLg it is \(f(x)=x-2\)

OpenStudy (anonymous):

no the first part is right but then after the and its f(g(-1))

OpenStudy (anonymous):

3

OpenStudy (anonymous):

\[1+7-9=-1\] i think

OpenStudy (anonymous):

no thats not an option lol this is hard the options are -21 -3 3 21

OpenStudy (anonymous):

then \[f(-1)=-1-2=-3\] for a final answer

OpenStudy (anonymous):

I got -3

OpenStudy (anonymous):

it looks hard because it takes two steps, not one first you find \(g(-1)\) then you find \(f\) of whatever you got for the first answer

OpenStudy (anonymous):

thank you!! im not sure who to give the medal too.